OpenStudy (buttermequeasy):

Given the function f(x) = 2(x + 4), find x if f(x) = 20. 6 48 14 8

OpenStudy (anonymous):

20=2x-8 28=2x x=14

OpenStudy (anonymous):

you are all wrong 20=2(x+4) 10=x+4 x=6

OpenStudy (anonymous):

2(x+4)=20 (x+4)=10 x=10-4 x=6

jimthompson5910 (jim_thompson5910):

f(x) = 2(x + 4) 20 = 2(x + 4) 20 = 2x + 8 20-8 = 2x 12 = 2x 12/2 = x 6 = x x = 6

OpenStudy (anonymous):

liamarie the mistake you made is u said -8 instead of +8 therefore, like i said, the answer is 6